OpenStudy (anonymous):

Show all work to simplify 2 over x + 2 over x plus 1 - 2 over x plus 2. Use complete sentences to explain how to simplify this expression. Remember to list all restrictions.

OpenStudy (anonymous):

@pengembara_bumi1

OpenStudy (anonymous):

@april115

OpenStudy (anonymous):

first..make sure the base are same..serenity...make ur base same..show me :)

OpenStudy (anonymous):

@serenity3 ???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

show me how u make the base same :)

OpenStudy (anonymous):

forget that question :/ help me factor this out please ? (y-1)(y+6)/(y-5)(y+2) x (y-5)(y+1)/(y-1)(y+3)

OpenStudy (anonymous):

\[\frac{ 2 }{ x } + \frac{ 2 }{ x+1 } - \frac{ 2 }{ x+2 }....\]

OpenStudy (anonymous):

ha ?? hahha..okok :)

OpenStudy (anonymous):

after that then we can keep working on the 2/x one D:

OpenStudy (anonymous):

make ur base be x+2

OpenStudy (anonymous):

for which one?